# An exact approach for the multi-depot electric vehicle scheduling problem

### v1

@@ Abstract
We propose a set of instances with up to 80 service trips to compare the two approaches, showing that, with a small number of depots, the compact 3-index model performs very well. However, as the number of depots increases the developed branch-and-cut algorithm proves to be of value. These findings not only offer algorithmic insights but the developed approaches also provide actionable guidance for transit agencies and operators, allowing to quantify trade-offs between fleet size, energy efficiency, and infrastructure needs under realistic operational conditions.
@@ 1 Introduction
In order to support the transformation of the public bus system in European countries and beyond, high-quality decision support is necessary, especially, concerning the bus operators’ important strategic decision of how many and which types of zero emission buses to invest in. Given an available line plan, switching from diesel-powered buses to zero emission vehicles without changes in the plan, requires solving an (electric) vehicle scheduling problem, minimizing the number of buses required to serve the line. Since battery-electric zero emission vehicles are currently still only available with rather limited driving ranges compared to diesel-powered vehicles, and have the disadvantage of taking a long time to recharge, it is not straightforward to substitute conventional diesel buses with zero emission technology, in particular for regional bus lines. Commercial transport, especially the schedule-based sector, operates with a high level of planning to ensure schedules are performed with high punctuality according to the timetable. Effectively integrating electric buses into these systems requires careful consideration of their operational constraints (Wen et al.,, 2016). Recharging stops during the day are usually required and have to be planned. From an operator’s perspective, charging on their own premises, potentially relying on their own photovoltaic energy, is in many cases the cheapest and sometimes the only option, since chargers for heavy duty vehicles are not yet available on a larger scale, resulting in a rather small number of possible charging locations.
@@
The remainder of this paper is structured as follows. We first give an overview of the work related to the problem setting considered in this paper in Section 2. We then formally define the MDEVSP with partial recharging and detail a graph representation of it in Section 3. Next, we present two mathematical formulations for the MDEVSP, one with three and one with two indices in Section 4. We propose a branch-and-cut algorithm to solve this problem in Section 5. The algorithm is tested on newly generated MDEVSP benchmark instances based on instances from the literature. Computational results are provided in Section 6, followed by the conclusions in Section 7.
@@ 2 Literature review
Electric vehicle scheduling problems have attracted quite some attention over the past years. In the following, we review some of the most closely related contributions, starting with the work of Li, (2014) who studied the single-depot electric vehicle scheduling problem (EVSP) without partial charging, and have shown that it is NP-hard. They develop a branch-and-price approach, considering battery swapping or fast charging up to battery capacity in conjunction with maximum distance constraints.
Wen et al., (2016) present a MILP for the MDEVSP considering partial charging relying on an almost acyclic network representation but requiring time variables, minimizing a combination of travel and vehicle costs. Larger instances are addressed by a large neighborhood search algorithm.
@@
Wang et al., (2021) develop a column generation algorithm and combine it with a genetic algorithm to address the MDEVSP of three bus lines in Qingdao, China. Adler and Mirchandani, (2017) address the the MDEVSP with partial recharge, minimizing the total schedule costs, limiting the number of buses, which can be stationed at each depot. A branch-and-price algorithm is developed and used to benchmark a heuristic approach on random instances with up to 50 service trips. The heuristic is then applied to a large-scale data set from the metropolitan area of Phoenix. Janovec and Koháni, (2019) develop a MILP for the single-depot problem with electric vehicles, considering possible charging events at each of the available chargers. Their formulation requires 4-index variables and, similar to our approach, they do not rely on time variables explicitly. The model is applied to case study-based data and solved by Xpress IVE. Up to 160 service trips are considered.
@@
Hu et al., (2022) address the optimization of locating fast chargers, which allow en-route charging at selected bus stops along three bus routes in Sydney, and determining charging schedules assuming time-dependent electricity prices. They also consider the possibility to delay service, but penalize passenger waiting times. Passenger demand as well as travel times are assumed to be uncertain. The developed robust optimization model is solved by Gurobi. Time-of-use electricity prices are also considered by van Kooten Niekerk et al., (2017), Li et al., (2020), and Wu et al., (2022) in a deterministic setting. van Kooten Niekerk et al., (2017) present two mathematical programming models for the single-depot case, where in the first, a linear charging process with constant electricity prices during the day is assumed. The second model allows any type of charging process, includes time-of-use electricity prices, and takes the depreciation cost of the battery into account. For the latter model, the exact value charged is approximated by discretizing it. Instances with up to 175 service trips are solved to optimality with both models, for larger instances with 241 service trips, methods based on column generation are deployed. Wu et al., (2022) formulate the problem with two objectives, which are considered in a lexicographic fashion and develop a branch-and-price algorithm. Also Liu and Ceder, (2020) address the combined problem of vehicle scheduling and charger location, allowing partial recharge for transit buses in Shanghai. The developed modeling approaches rely on deficit function theory and mixed-integer programming. The location of charging infrastructure in combination with electric bus scheduling and a single depot is modeled as a 2-index MILP by Stumpe, (2024) and, using Gurobi as the solver engine, applied to large-scale instances. Zhou et al., (2024) propose a MILP and a set-covering formulation for the electric bus charging scheduling problem with a mixed fleet of electric vehicles, partial recharging while taking battery degradation effects into account. A branch-and-price algorithm solves instances with up to 100 service trips, for tackling large-scale instances, an optimization-based adaptive large neighborhood search (ALNS) is developed.
@@
A mixed fleet of conventional and electric vehicles is also considered by Sassi and Oulamara, (2017). Since charging costs change over time, in the developed model, time is discretized and costs are minimized in the objective function. Also, maximum grid capacity constraints are considered. Rinaldi et al., (2020) address mixed fleet bus scheduling with a single depot. Time is discretized, assuming that buses can be fully charged within one time interval. Trip departure times may deviate from their preferred time. The developed MILP is decomposed into smaller subproblems and applied to case study data from Luxembourg City. Yıldırım and Yıldız, (2021) also address a mixed fleet problem, but with multiple depots and with multiple different charging technologies and develop an efficient column generation-based algorithm to determine the minimum cost fleet configuration for large-scale real-world instances. Zhang et al., (2022) present a mixed-integer programming formulation for the MDEVSP with heterogeneous vehicles. They solve the model using CPLEX and propose an ALNS algorithm to address the problem, incorporating a partial mixed-route strategy and a partial recharging policy. Frieß and Pferschy, (2024) have developed a MILP, where different zero emission propulsion technologies are considered concurrently to optimize the bus fleet mix for serving the city of Graz, Austria. In order to solve the model, they pre-select a set of possible options for buses to switch between different lines.
@@
Zhang et al., (2021) address electric bus fleet scheduling with a single depot, non-linear charging and battery fading, and develop a branch-and-price approach. Non-linear charging, considering battery degradation effects are also studied by Zhou et al., (2022). A column-generation based heuristic enhanced by ideas from machine learning has recently been developed by Gerbaux et al., (2025) and applied to the MDEVSP with non-linear charging. Also Diefenbach et al., (2022) consider non-linear charging in an in-plant vehicle scheduling application of the multi-depot problem. In their case, vehicles are free to return to any depot, not necessarily the one they started from. They address this problem setting by a branch-and-check approach, moving all complicating aspects, such as the planning of charging events, to the subproblem and generate cutting planes, enforcing a change in the current solution, when violated. Löbel et al., (2024) propose a MILP for the electric bus scheduling problem with a single depot for mixed fleets of electric and non-electric vehicles, presenting an improved approximation of non-linear battery charging behavior of the electric vehicles. Furthermore, they address the challenge of adjusting to power grid bottlenecks by integrating dynamic recharging rates and time-of-use electricity prices. Partial charging is allowed and available charging slots must not be exceeded. The model is applied on diverse real-life instances with up to 1,207 service trips.
Capacitated charging stations in conjunction with partial recharging in the context of electric vehicle scheduling in public transit is considered by de Vos et al., (2024). They rely on a path-based formulation which is solved by column generation. Integer solutions are obtained by price-and-branch and a diving heuristic. Instances with up to 816 service trips are addressed.
@@ 3.1 Description of the MDEVSP
Furthermore, the maximum battery capacity of the electric vehicles $s^{max}$, the minimum allowed battery level $s^{min}$, the minimum allowed battery level when returning to the depot at the end of the schedule $s^{min}_{dep}$ (with $s^{min}_{dep}\geq s^{min}$), the minimum time that needs to be available for charging $t^{min}$ and the charging rate $r$, i.e., the time needed to charge one unit of the battery, are given. From that we can compute the time required to fully charge the battery $t^{max}$ as $t^{max}=\frac{1}{r}(s^{max}-s^{min})$. A vehicle’s charging time is assumed to be a linear function of the amount that its battery is charged.
The goal of the MDEVSP is to find a schedule, such that (1) each timetabled service trip $i\in V^{I}$ is done by exactly one vehicle, (2) each vehicle starts in some depot $k\in K$, performs a sequence of service trips and returns to the same depot $k$ at the end of the day, (3) between two service trips or between a service trip and returning to the depot, each vehicle visiting a charging station can reload its battery for at least $t^{min}$ minutes with charging rate $r$, (4) each vehicle starts at its depot with energy level $s^{max}$ and returns to its depot with energy level at least $s^{min}_{dep}$ at the end of the day, (5) the energy level of each vehicle never drops below $s^{min}$, and (6) in each depot $k\in K$ at most $b_{k}$ vehicles start and end their routes. We call such a schedule feasible. Among all feasible schedules the MDEVSP searches for the one which lexicographically minimizes (i) the number of used vehicles, (ii) the number of visits of charging stations, and (iii) the energy spent on deadhead trips. The latter is is equivalent to minimizing the total energy consumed since the energy spent on service trips is a constant value.
@@ 3.2 Graph representation
    Generate an arc between each pair of time-feasible service trips: For each pair of service trips $i,j\in V^{I}$, we create an arc if $s_{i}+u_{i}+t_{ij}\leq s_{j}$ holds.
@@
    Generate charging nodes and connect (full) charging nodes representing charging between two service trips: For each pair of service trips $i,j\in V^{I}$, we identify the set $C_{ij}\subseteq C$ of all charging stations, which are reasonable for charging a vehicle between the service trips $i$ and $j$. To be more precise, a charging station $a\in C$ is in $C_{ij}$, if there is no other charging station $a^{\prime}\in C$ that dominates $a$. A charging station $a^{\prime}$ dominates a charging station $a$, if (1) $a^{\prime}$ is closer to the end location $\ell^{e}_{i}$ of service trip $i$ than $a$ and (2) $a^{\prime}$ is closer to the start location $\ell^{s}_{j}$ of service trip $j$ than $a$ (if such a charging station $a^{\prime}$ exists, then it would always be better to charge at $a^{\prime}$ than at $a$).
@@
    On the one hand, if we can fully charge, so if $t_{ija}=t^{max}$, then we use the full charging node $c=c^{full}_{ia}$, we connect $c$ and $j$ with an arc and we set $p_{cj}=p_{aj}$.
@@
The travel time $t_{ij}$ between service trips, depots, and charging stations, as well as the start time $s_{i}$ and end time $e_{i}$ of each service trip can be found in Table 1. The time $t^{max}$ it takes to fully charge a vehicle at a full charging node $c^{full}_{ia}$ is set to $120$ minutes. Recharging at the partial charging nodes $c^{part}_{121}$ and $c^{part}_{122}$ is possible for $72$ and $116$ minutes, respectively. However, since charging station $a_{2}$ is closer to $ST_{1}$ as well as to $ST_{2}$ than $a_{1}$, $ST_{1}$ and $ST_{2}$ are only connected via $c^{part}_{122}$. The same holds for $ST_{1}$ and $ST_{3}$, where enough time is available to fully recharge, but again $a_{2}$ is closer to both $ST_{1}$ and $ST_{3}$ than $a_{1}$. Therefore, $ST_{1}$ and $ST_{3}$ are only connected via $c^{full}_{12}$. An example of the graph for instance $I_{1}$ is provided in Figure 1.
@@
Grey arcs are direct connections between service trips and between service trips and depots. Black dashed arcs connect service trips with charging nodes, as well as charging nodes with the depot. All connections are time feasible. E.g., $ST_{2}$ and $ST_{3}$ are not connected via an arc, since $ST_{2}$ ends at 5:15 p.m. but $ST_{3}$ already starts at 5:05 p.m. All service trips are connected via two full charging nodes to the end depot $d_{1}$, because charging station $a_{1}$ is closer to the depot while $a_{2}$ is closer to the service trips. Thus, connections via full charging nodes have been generated for all combinations of service trips and charging stations.

### v2

@@ Abstract
We propose a set of instances with up to 80 service trips to compare the two approaches, showing that, with a small number of depots, the compact 3-index model performs very well. However, as the number of depots increases, the developed branch-and-cut algorithm proves to be of value. The inclusion of realistic instances and technology-specific scenarios for diesel, battery-electric, and fuel cell-electric buses further strengthens the practical relevance of our results, offering concrete guidance for sustainable fleet planning.
Keywords: electric bus scheduling, electric vehicle scheduling problem, branch-and-cut, multiple depots, partial charging, integer programming
Acknowledgments
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ 1 Introduction
In order to support the transformation of the public bus system in European countries and beyond, high-quality decision support is necessary, especially, concerning the bus operators’ important strategic decision of how many and which types of zero emission buses to invest in. Given an available line plan, switching from diesel-powered buses to zero emission vehicles without changes in the plan, requires solving an (electric) vehicle scheduling problem, minimizing the number of buses required to serve the line. Since battery-electric zero emission vehicles are currently still only available with rather limited driving ranges compared to diesel-powered vehicles, and have the disadvantage of taking a long time to recharge, it is not straightforward to substitute conventional diesel buses with zero emission technology, in particular for regional bus lines. Recharging stops during the day are usually required and have to be planned. From an operator’s perspective, charging on their own premises, potentially relying on their own photovoltaic energy, is in many cases the cheapest and sometimes the only option, since chargers for heavy duty vehicles are not yet available on a larger scale, resulting in a rather small number of possible charging locations.
@@
  • We compare propulsion technologies (diesel, battery-electric, and fuel cell-electric) using realistic instances, conducting a parameter study to examine the impact of cold weather conditions as well as battery degradation on fleet size and operational feasibility.
The remainder of this paper is structured as follows. We first give an overview of the work related to the problem setting considered in this paper in Section 2. We then formally define the MDEVSP with partial recharging and detail a graph representation of it in Section 3. Next, we present two mathematical formulations for the MDEVSP, one with three and one with two indices in Section 4. We propose a branch-and-cut algorithm to solve this problem in Section 5. The algorithm is tested on newly generated MDEVSP benchmark instances based on an instance generation scheme commonly used in the literature. Computational results are provided in Section 6, followed by a comparison of propulsion technologies in Section 7. Conclusions are drawn in Section 8.
@@ 2 Literature review
Electric vehicle scheduling problems have attracted quite some attention over the past years. In the following, we review some of the most closely related contributions, starting with the work of Li, (2014) who studied the single-depot electric vehicle scheduling problem (EVSP) without partial charging, and have shown that it is NP-hard. They develop a branch-and-price approach, considering battery swapping or fast charging up to battery capacity in conjunction with maximum distance constraints. Capacitated charging stations in conjunction with partial recharging in the context of electric vehicle scheduling in public transit is considered by de Vos et al., (2024). They rely on a path-based formulation which is solved by column generation. Integer solutions are obtained by price-and-branch and a diving heuristic. Instances with up to 816 service trips are addressed.
@@
Janovec and Koháni, (2019) develop a MILP for the single-depot problem with electric vehicles, considering possible charging events at each of the available chargers. Their formulation requires 4-index variables and, similar to our approach, they do not rely on time variables explicitly. The model is applied to case study-based data with up to 160 service trips and solved by Xpress IVE.
Wen et al., (2016) present a MILP for the MDEVSP considering partial charging relying on an almost acyclic network representation but requiring time variables, minimizing a combination of travel and vehicle costs. Larger instances are addressed by a large neighborhood search algorithm. Wang et al., (2021) develop a column generation algorithm and combine it with a genetic algorithm to address the MDEVSP of three bus lines in Qingdao, China. Adler and Mirchandani, (2017) address the the MDEVSP with partial recharge, minimizing the total schedule costs, limiting the number of buses, which can be stationed at each depot. A branch-and-price algorithm is developed and used to benchmark a heuristic approach on random instances with up to 50 service trips. The heuristic is then applied to a large-scale data set from the metropolitan area of Phoenix.
@@
Hu et al., (2022) address the optimization of locating fast chargers, which allow en-route charging at selected bus stops along three bus routes in Sydney, and determining charging schedules assuming time-dependent electricity prices. Further they consider the possibility to delay service, but penalize passenger waiting times. Passenger demand as well as travel times are assumed to be uncertain. The developed robust optimization model is solved by Gurobi. Time-of-use electricity prices are also considered by van Kooten Niekerk et al., (2017), Li et al., (2020), and Wu et al., (2022) in a deterministic setting. van Kooten Niekerk et al., (2017) present two mathematical programming models for the single-depot case, where in the first, a linear charging process with constant electricity prices during the day is assumed. The second model allows any type of charging process, includes time-of-use electricity prices, and takes the depreciation cost of the battery into account. For the latter model, the exact value charged is approximated by discretizing it. Instances with up to 175 service trips are solved to optimality with both models, for larger instances with 241 service trips, methods based on column generation are deployed. Wu et al., (2022) formulate the problem with two objectives, which are considered in a lexicographic fashion and develop a branch-and-price algorithm. Also Liu and Ceder, (2020) address the combined problem of vehicle scheduling and charger location, allowing partial recharge for transit buses in Shanghai. The developed modeling approaches rely on deficit function theory and mixed-integer programming. The location of charging infrastructure in combination with electric bus scheduling and a single depot is modeled as a 2-index MILP by Stumpe, (2024) and, using Gurobi as the solver engine, applied to large-scale instances.
@@
Zhou et al., (2024) propose a MILP and a set-covering formulation for the electric bus charging scheduling problem with a mixed fleet of electric vehicles, partial recharging while taking battery degradation effects into account. A branch-and-price algorithm solves instances with up to 100 service trips, for tackling large-scale instances, an optimization-based adaptive large neighborhood search (ALNS) is developed. A mixed fleet of conventional and electric vehicles is also considered by Sassi and Oulamara, (2017). Since charging costs change over time, in the developed model, time is discretized and costs are minimized in the objective function. Also, maximum grid capacity constraints are considered. Rinaldi et al., (2020) address mixed fleet bus scheduling with a single depot. Time is discretized, assuming that buses can be fully charged within one time interval. Trip departure times may deviate from their preferred time. The developed MILP is decomposed into smaller subproblems and applied to case study data from Luxembourg City. Yıldırım and Yıldız, (2021) also address a mixed fleet problem, but with multiple depots and with multiple different charging technologies and develop an efficient column generation-based algorithm to determine the minimum cost fleet configuration for large-scale real-world instances. Zhang et al., (2022) present a mixed-integer programming formulation for the MDEVSP with heterogeneous vehicles. They solve the model using CPLEX and propose an ALNS algorithm to address the problem, incorporating a partial mixed-route strategy and a partial recharging policy. Frieß and Pferschy, (2024) have developed a MILP, where different zero emission propulsion technologies are considered concurrently to optimize the bus fleet mix for serving the city of Graz, Austria. In order to solve the model, they pre-select a set of possible options for buses to switch between different lines.
@@
Zhang et al., (2021) address electric bus fleet scheduling with a single depot, non-linear charging and battery fading, and develop a branch-and-price approach. Non-linear charging, considering battery degradation effects are also studied by Zhou et al., (2022). A column-generation based heuristic enhanced by ideas from machine learning has recently been developed by Gerbaux et al., (2025) and applied to the MDEVSP with non-linear charging. Also Diefenbach et al., (2022) consider non-linear charging in an in-plant vehicle scheduling application of the multi-depot problem. In their case, vehicles are free to return to any depot, not necessarily the one they started from. They address this problem setting by a branch-and-check approach, moving all complicating aspects, such as the planning of charging events, to the subproblem and generate cutting planes, enforcing a change in the current solution, when violated. Löbel et al., (2024) propose a MILP for the electric bus scheduling problem with a single depot for mixed fleets of electric and non-electric vehicles, presenting an improved approximation of non-linear battery charging behavior of the electric vehicles. Furthermore, they address the challenge of adjusting to power grid bottlenecks by integrating dynamic recharging rates and time-of-use electricity prices. Partial charging is allowed and available charging slots must not be exceeded. The model is applied to diverse real-life instances with up to 1,207 service trips.
@@ 3.1 Description of the MDEVSP
Furthermore, the maximum battery capacity of the electric vehicles $s^{max}$, the minimum allowed battery level $s^{min}$, the minimum allowed battery level when returning to the depot at the end of the schedule $s^{min}_{dep}$ (with $s^{min}_{dep}\geq s^{min}$), the minimum time that needs to be available for charging $t^{min}$ and the charging rate $r$, i.e., the time needed to charge one unit of the battery, are given. From that we can compute the time required to fully charge the battery $t^{max}$ as $t^{max}=(s^{max}-s^{min})/r$. A vehicle’s charging time is assumed to be a linear function of the amount that its battery is charged.
The goal of the MDEVSP is to find a schedule, such that (1) each timetabled service trip $i\in V^{I}$ is done by exactly one vehicle, (2) each vehicle starts in some depot $k\in K$, performs a sequence of service trips and returns to the same depot $k$ at the end of the day, (3) between two service trips or between a service trip and returning to the depot, each vehicle visiting a charging station can reload its battery for at least $t^{min}$ minutes with charging rate $r$, (4) each vehicle starts at its depot with energy level $s^{max}$ and returns to its depot with energy level at least $s^{min}_{dep}$ at the end of the day, (5) the energy level of each vehicle never drops below $s^{min}$, and (6) in each depot $k\in K$ at most $b_{k}$ vehicles start and end their routes. We call such a schedule feasible. Among all feasible schedules the MDEVSP searches for the one which lexicographically minimizes (i) the number of used vehicles, (ii) the number of visits of charging stations, and (iii) the energy spent on deadhead trips. The latter is equivalent to minimizing the total energy consumed, as the energy spent on service trips is a constant value.
@@ 3.2 Graph representation
    Generate an arc between each pair of time-feasible service trips: For each pair of service trips $i,j\in V^{I}$, we create an arc if $s_{i}+u_{i}+t_{ij}\leq s_{j}$ holds. We store these pairs of time-feasible service trips in $F$.
@@
    Generate charging nodes and connect (full) charging nodes representing charging between two service trips: For each pair of service trips $(i,j)\in F$, we identify the set $C_{ij}\subseteq C$ of all charging stations, which are reasonable for charging a vehicle between the service trips $i$ and $j$. To be more precise, a charging station $a\in C$ is in $C_{ij}$, if there is no other charging station $a^{\prime}\in C$ that dominates $a$. A charging station $a^{\prime}$ dominates a charging station $a$, if (1) $a^{\prime}$ is closer to the end location $\ell^{e}_{i}$ of service trip $i$ than $a$ and (2) $a^{\prime}$ is closer to the start location $\ell^{s}_{j}$ of service trip $j$ than $a$ (if such a charging station $a^{\prime}$ exists, then it would always be better to charge at $a^{\prime}$ than at $a$).
@@
    On the one hand, if we can fully charge, so if $t_{ija}=t^{max}$, then we use the full charging node $c=c^{full}_{ia}$. We connect $c$ and $j$ with an arc and we set $p_{cj}=p_{aj}$.
@@
The travel time $t_{ij}$ between service trips, depots, and charging stations, as well as the start time $s_{i}$ and end time $e_{i}$ of each service trip can be found in Table 1. The time $t^{max}$ it takes to fully charge a vehicle at a full charging node $c^{full}_{ia}$ is set to $120$ minutes. Recharging at the partial charging nodes $c^{part}_{121}$ and $c^{part}_{122}$ is possible for $72$ and $116$ minutes, respectively. However, since charging station $a_{2}$ is closer to both $ST_{1}$ and $ST_{2}$ than $a_{1}$, $ST_{1}$ and $ST_{2}$ are only connected via $c^{part}_{122}$. The same holds for $ST_{1}$ and $ST_{3}$, where enough time is available to fully recharge, but again $a_{2}$ is closer to both $ST_{1}$ and $ST_{3}$ than $a_{1}$. Therefore, $ST_{1}$ and $ST_{3}$ are only connected via $c^{full}_{12}$. An example of the graph for instance $I_{1}$ is provided in Figure 1.
@@
Grey arcs are direct connections between service trips and between service trips and depots. Black dashed arcs connect service trips with charging nodes, as well as charging nodes with the depot. All connections are time feasible. E.g., $ST_{2}$ and $ST_{3}$ are not connected via an arc, since $ST_{2}$ ends at 5:15 p.m. but $ST_{3}$ already starts at 5:05 p.m. All service trips are connected via two full charging nodes with the destination depot $d_{1}$, since charging station $a_{1}$ is closer to the depot while $a_{2}$ is closer to the service trips. Thus, connections via full charging nodes have been generated for all combinations of service trips and charging stations.